# biblio.bib

### v2

@@ A.1 Obtaining the optimal solution in the memoryless case
Previous work [fischer2024minimising] has shown that the optimal pulsing protocol for the memoryless case requires an initial drug application phase until the resistant fraction reaches some upper threshold $\phi_{h}$, followed by a pause phase in which the resistant fraction decreases to some lower bound $\phi_{l}$. This is followed by repeated cycles of drug treatment and pause phases where the switching time occurs when the resistant fraction reaches $\phi_{h}$ and $\phi_{l}$, respectively. To obtain the optimal values of $\phi_{h}$ and $\phi_{l}$ for our specific model parameterization, we swept over values of $\phi_{h}$ and $\phi_{l}$ between 0.1 and 0.9, with increments of 0.04, selecting the values which yielded the highest net death rate ($\phi_{l}=0.48$ and $\phi_{h}=0.52$).
Appendix B Reinforcement Learning Details
@@
We adapted DQN from Stable-Baselines3 [stable-baselines3] with a Double DQN action selection rule [van2016deep]. We train the RL agent for $3\times 10^{5}$ total environment steps. We limited the episodes to be of length $10^{4}$ steps (corresponding to 100 hours in simulation). An MLP of fixed size ($2$ hidden layers with $64$ dimensions each and ReLU activation) was used to parameterize the $Q$-function.
@@
Regarding the environment and training, we list several key implementation choices: At the beginning of an episode, the state is zero-padded to always be of length $K=5$. We experimented with adding a penalty for allowing the number of cells to increase beyond the initial amount upon termination, but found this was not necessary for successful training. Borrowing terminology from the literature on Atari environments [nature-dqn], we stack $K$ frames (previous cost values) to form the RL agent’s state vector. Although initially we let $K$ be a $\mu$-dependent hyperparameter, we found a constant choice of $K=5$ to work well across the values of $\mu$ studied.
@@
B.1 Hyperparameters
We find that sweeping over a range of hyperparameters (as shown in Table B.1) did not have a significant effect on performance, though for reproducibility we list the final hyperparameters used (for $\mu=0.7$) below in Table B.1.
Table 1: Finetuned Hyperparameter Values for Double DQN
Hyperparameter Finetuned Value
batch size $32$
buffer size $100{,}000$
exploration rate $0.05$
frames stacked $5$
gradient steps (UTD / RR) $1$
learning rate $3.60\times 10^{-4}$
target update interval $1,000$
discount factor $0.999$
learning starts $10,000$
Table 2: Hyperparameter Sweep Ranges
Hyperparameter Sweep Values
batch size $16$, $32$, $64$
exploration rate $0.01-0.2$
learning rate $10^{-5}-10^{-3}$
target update interval $1{,}000,5{,}000,10{,}000,30{,}000$
target Polyak averaging $0.95,0.99,0.995,1.0$
Appendix C Further Experimental Results
Figure 3: The learned policy shows a resemblance to the optimal memoryless strategy, with an initial constant application phase followed by a pulsatile phase. However, in the case of memory-based dynamics, the frequency of pulsing must be increased over time as discussed in Sec. 3. Since the policy is eventually limited by the simulation time, the pulsing frequency becomes bottlenecked by our choice of time discretization $\Delta$ after $\approx 20$ hours. Despite this, the policy is still able to perform well with rapid pulsing.
Figure 4: We find that the simulation time can have a significant effect on RL performance. For each choice of $\Delta$, we run a random sweep of size $30$ over various hyperparameters, selecting the highest-performing run for each $\Delta$. Since smaller values of $\Delta$ require longer compute-times for simulations, there is a tradeoff between the amount of time (also, the inverse of max pulsing rate, which may be more relevant in clinical settings) and the best performance. We have chosen to use $\Delta=0.01$ throughout.

### v3

@@ 2.4 Obtaining the optimal solution in the memoryless case
As mentioned above, previous work (Fischer & Bluethgen, 2024) showed that the optimal pulsing protocol for the memoryless case requires an initial drug application phase until the resistant fraction reaches some upper threshold $\phi_{h}$, followed by a pause phase (“drug holiday”) in which the resistant fraction naturally relaxes back to some lower bound $\phi_{l}$. This is followed by similarly repeated cycles of treatment and pause phases where the switching time occurs when the resistant fraction reaches $\phi_{h}$ and $\phi_{l}$, respectively. To obtain the optimal values of $\phi_{h}$ and $\phi_{l}$ for our specific model parameterization, we swept over values of $\phi_{h}$ and $\phi_{l}$ between $0.1$ and $0.9$, with increments of $0.011$, selecting the values which yielded the highest net death rate: $\phi_{l}=0.494$ and $\phi_{h}=0.505$.
To compare as baselines against our learned policy in non-Markovian environments, we repeated the same procedure for different values of $\mu$. The optimal switching fractions are given in Table 1.
$\mu$ $\phi_{l}$ $\phi_{h}$
$1.0$ $0.494$ $0.505$
$0.9$ $0.483$ $0.494$
$0.8$ $0.466$ $0.477$
$0.7$ $0.448$ $0.459$
$0.6$ $0.428$ $0.439$
Table 1: Optimal switching fraction values for the baseline policy for each memory strength $\mu$.
3 Reinforcement Learning
Refer to caption Refer to caption
Figure 2: Left: The learned policy shows a resemblance to the optimal memoryless strategy, with an initial constant application phase followed by a pulsatile phase. However, in the case of memory-based dynamics, the frequency of pulsing must be increased over time as discussed in Sec. 4. Since the policy is eventually limited by the simulation time, the pulsing frequency becomes bottlenecked by our choice of time discretization $\Delta$ after $\approx 20$ hours. Despite this, the policy is still able to perform well with rapid pulsing. Right: Effect of learned policy on resistant fraction. for different memory strengths. The RL agent finds (for distinct $\mu$ values) appropriate lower and upper bounds for the fraction of resistant cells. Maintaining the subpopulation in this range ensures the population can be controlled.
Reinforcement learning (RL) allows an agent to learn an optimal control strategy (policy) directly from the experience it collects. This data may be from simulation (Akkaya et al., 2019), or from real-world data (Haarnoja et al., 2024), or even from fully offline datasets (Levine et al., 2020). In RL, the interactive agent receives observations from the environment. As a result of this input, the agent outputs a control or action (denoted $u$ throughout), yielding a transition to a new state and a scalar reward (negative cost). The goal of the agent, analogous to OC, is to act in a way that maximizes the expected long-term accumulated rewards. In our problem formulation, this will correspond to the use of drug to minimize the total population count.
We consider the online RL setting in simulation, where a drug protocol is learned through experience (Fig. 2, left) by continually improving a policy as it is learned. Despite the non-Markovian dynamics and even without access to the underlying environment-specific model parameters, we show that RL is capable of providing high-performing policies. To formulate the RL problem, we define the relevant characteristics as follows (in the following, let $\Delta$ denote the simulation time between actions):
State: The state ($s_{t}$) of the agent’s environment is a list of the last $K=5$ estimates of the instantaneous growth rate and one-hot encoded actions, $s_{t}=[c_{t-K,..t},u_{t-K,..t}]$ where ${c_{t}={\Delta^{-1}}\log N_{t}/N_{t-\Delta}}$ is a pointwise version of the cost defined above. Thus, a history of past observations is encoded in the state vector, a crucial design choice if the agent is to learn control without a recurrent hidden state. Action: As motivated in Section 2.1, we choose a binary action space $u\in\{0,1\}$ representing whether the drug is applied or not, as we have proven that this is sufficient to recover optimal control. Reward: As we seek to solve Problem 4, the reward is simply the negative growth rate, $r_{t}=-c_{t}$. Notice that with this choice of reward function, the sum of rewards across a trajectory simplifies as $R_{0:T}=\Delta^{-1}\log N_{0}/N_{T}$, ensuring the agent’s objective is aligned with a reduction in total cell population. Dynamics: The dynamics of the total cell population $N_{t}$ are governed by the dynamical system described in Section 2.1, initialized to be fully susceptible ($\textbf{x}_{0}=[1000,0]$). After an action is executed, the simulation (a numerical solution  [^1] The numerical solution of fractional differential equations requires some care; cf. (Garrappa, 2018) for details. of equation 3) is computed with the action (dose) fixed for $\Delta=0.01$ hours to compute the next state ($s_{t+\Delta}$). To encourage the agent to reduce the cell population while decreasing computational runtime and allowing for exploration, we terminate the episode if the number of cells ever exceeds $150\%$ of its initial amount.
The goal in RL is to learn the policy $\pi^{*}$ which maximizes the expected cumulative discounted sum of rewards across a trajectory:
$\pi^{*}(a|s)=\operatorname*{arg\,max}_{\pi}\mathbb{E}_{\tau\sim{}p,\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}r(s_{t},u_{t})\right].$ (12)
To solve this optimization problem, we consider discrete-action value-based methods. Within this framework, the optimal action-value function $Q^{*}(s,u)$ is learned based on off-policy data observed during exploration. The value function satisfies the Bellman optimality equation (written to match the control notation introduced above):
Figure 3: Performance comparison of constant drug application, solution for the memoryless case, resistant fraction-based pulsing technique, and policy learned by RL. For the fraction-based policy, an optimal lower and upper bound for resistant fractions are found through sweeping (Appendix 2.4). The RL policy is capable of controlling the cell population better than any other scheme.
Within this class of methods, DQN (Mnih et al., 2015) learns $Q^{*}$ by exploring with an annealed greedy exploration strategy. After an action is taken, the experience tuple $(s,a,r,s^{\prime})$ is stored in an experience replay buffer for later use. In deep RL, the action-value function is parameterized with a neural network (denoted by the trainable parameters $\theta$), and trained via SGD by sampling mini-batches, denoted $\mathcal{B}$, uniformly at random from the buffer. The temporal-difference loss function is defined as the Bellman residual – the squared difference between left and right hand sides of equation 13:
$\mathcal{L}_{\theta}=\sum_{\{s,u,r,s^{\prime}\}\in\mathcal{B}}\left|Q^{*}(s_{t},u_{t};\mathbf{\theta})-\left(r+\gamma\max_{u^{\prime}}Q^{*}(s_{t+\Delta},u^{\prime};\mathbf{\bar{\theta}})\right)\right|^{2}\;.$ (14)
As common in value-based algorithms, we use an additional target network to reduce the effect of bootstrapping during training. The target network (denoted with parameters $\bar{\theta}$) has its parameters frozen during training, and updated periodically directly by copying the online network parameters $\theta\to\bar{\theta}$. We additionally use the double DQN approach to mitigate the over-estimation bias in standard DQN (Van Hasselt et al., 2016). Further details on reinforcement learning can be found e.g. in (Sutton & Barto, 2018; Hessel et al., 2018). We tune over several hyperparameters (whose values we list in the Appendix) to optimize performance. Our implementation of double DQN (Van Hasselt et al., 2016) (and later FQF) are based on open-source code from Stable-Baselines3 (Raffin et al., 2021). Each agent is trained for $3\times 10^{5}$ environment steps. All code to reproduce our experimental results can be found at https://github.com/JacobHA/RL4Dosing.
4 Case Study of Fixed Memory Environments
We first test DQN in the memoryless case (${\mu=1}$), for which an optimal controller is known (Fischer & Bluethgen, 2024). In this case, the optimal policy can be derived based only on two consecutive resistant fractions. However, recall that the RL agent is only given access to the instantaneous growth rates, not the resistant fractions. Nevertheless, the RL agent is surprisingly able to reliably recover the optimal policy with only two successive frames and without any access to underlying model parameters. The control strategy found by the agent involves first applying the drug and then pulsing regularly, in agreement with the OC solution, as seen in Fig. 3 (rightmost plot). Indeed, Fig. 2 (right panel) shows that the agent discovers an internal representation based on growth rates that is compatible with fraction-based switching.
With confirmation that RL can find the optimal dosing strategy in the memoryless case, we turn to the more difficult memory-based dynamics ($\mu<1$). We do not have a solution to Problem 4 in this regime, so we compare against two baselines: The memoryless protocol described above and a modified version in which switching times occur when the resistant fraction reaches a threshold value. We note that these baselines can have arbitrarily small switching times, which is practically infeasible. Remarkably, we find that using a small but bounded $\Delta$, our learned policy outperforms both of these baselines (Fig. 3). In addition, our experiments show that decreasing $\Delta$ (increasing measurement frequency) beyond a certain threshold does not considerably increase performance (Fig. 6).
Interestingly, the learned policy reveals that the agent initially maintains constant drug application before transitioning to a memory-dependent pulsing protocol. The agent increases the frequency of pulsing throughout the trajectory until saturation at the maximum rate ($\Delta^{-1}$) (Fig. 2). For cases with memory, this increase in dosing frequency can be understood as the result of faster cellular adaptation back to the resistant state after multiple drug encounters. Thus, to maintain a susceptible population, the policy’s pulse frequency must continuously be increased to compensate for the memory-based adaptability. We also find that the agent maintains a lower average resistant fraction at higher memory strengths (Fig. 2, right). This qualitative trend is in agreement with the optimal switching fractions found by optimal control shown in Table 1.
Figure 4: PPO and SAC fail to find a bang-bang control policy and have a lower performance than DQN; highlighting the need for discrete action algorithms, as informed by optimal control.
To independently show the benefit of using discrete actions, the original problem can instead be approached with continuous action algorithms, such as PPO (Schulman et al., 2017) or SAC (Haarnoja et al., 2018). As shown in Fig. 4, these algorithms considerably underperform relative to DQN.
Qualitatively, we find that these continuous action algorithms successfully find the constant application period at the beginning of the episode. As the episode continues, though, both PPO and SAC attempt a short-lived bang-bang protocol before defaulting to the average action for the remainder of the episode. These results were obtained from an equal-compute hyperparameter sweep in a fixed $\mu=0.9$ environment. We conjecture that continuous action algorithms struggle because of the saturating gradient effect at the limits of action space, which correspond to bang-bang control. Additionally, continuous action algorithms that give bonuses for policy entropy (e.g. SAC) make the task of learning these deterministic policies even more difficult.
5 Model successfully generalizes to memory and observation noise perturbations
In the previous sections, we construct RL agents that are each trained on specific dynamics determined by a fixed $\mu$ value. However, this can be problematic in practice, as the memory strength $\mu$ is difficult to measure (it would require fitting equation 5 to entire episodes of data), which could correspond to adverse costs in the clinical setting. Furthermore, adaptive cell populations such as cancer can dynamically change their memory capacity over time (Ringrose & Paro, 2004; Acar et al., 2005). Thus, a useful policy should be able to interact immediately without requiring additional samples, and with no access to the value of $\mu$. The unknown value of $\mu$ thus represents “side information” that is not given to the agent but provides a context for the agent to understand the current environment. This places the new problem setting in the framework of contextual MDPs (Hallak et al., 2015; Sodhani et al., 2022). To approach this problem while maintaining clinical applicability, we train a general agent over a range of memory strengths. This approach of “domain randomization” (DR) has been applied with great success in other RL problems (Tobin et al., 2017; Akkaya et al., 2019).
To tackle this more challenging environment, we use FQF (Yang et al., 2019), a recent algorithm for distributional RL, where the distribution (rather than only the mean) of future returns is learned. FQF and its variants have shown success in discrete control environments. In particular, we use a noisy variant that combines the approach of NoisyNets (Fortunato et al., 2018) for improved exploration. Further, we find it useful to supplement the RL state with the one-hot encoded actions used in the last $K$ steps. This allows the agent to observe the response of the environment to the presence of drug: At lower $\mu$, application of drug has a drastically different effect than at higher $\mu$ values at similar growth rates. With these changes, we sweep over hyperparameters (cf. Appendix) to find a model capable of reducing cell populations given any memory strength.
Figure 5: Left: Policy is robust to observation noise, as tested in a memoryless environment. To represent potential measurement errors in the clinical setting, noise is drawn from a normal distribution with standard deviation $\sigma$ and added to each state before given to the agent. Despite large amounts of observation noise, the agent was able to drive population reduction and maintain a similar resistant fraction. Each trace represents the mean over $10$ trajectories, with the standard error represented by the shaded region. Right: Learned policy is general and robust to changes in memory strength. Every $20$ decision steps the memory strength is reset to a new value, drawn from a uniform distribution over the interval $[0.6,1]$. Agent is able to quickly adapt to mitigate population growth in the new environment.
Although the $\mu$-specific models outperform the generalist agent in their own finetuned environments, the generalist agent can achieve good performance across a range of memory strengths (Fig. 5). It is worth noting that this model now enables use in settings where memory strength is entirely unknown and possibly dynamical or noisy in nature. Experimentally, we find this agent is very robust to perturbations of various types: namely large changes in $\mu$ and large amounts of observation noise. In both cases, the learned policy employed by the agent is able to successfully adapt and continue to minimize population growth (Fig. 5).
6 Related Work
The use of reinforcement learning in control of biological systems, especially in dosage control, has been investigated by others. We will briefly highlight the most relevant prior work in this section. (Padmanabhan et al., 2017) studies a deterministic memory-less model of cancer dynamics with additional “immune” and “normal” cell types included. The number of tumor cells is used as a discrete RL state, and hence a tabular method is used. In (Engelhardt, 2020), a coupled SDE governs dynamics of $d$-many phenotypic subpopulations and the RL state is composed of measurements of individual phenotype populations. Although the specific dynamical model varies, prior work often requires unrealistic inputs to define the RL observation. An important contribution of the current work is in advancing the model class itself while maintaining clinical relevance by using easily measured quantities and providing a policy that generalizes across memory strengths.
7 Future Work
We utilize frame-stacking (Mnih et al., 2015) to encapsulate the history of the agent’s trajectory, but in future work we will test the use of recurrent policies to capture more nuanced long-term effects which may further improve performance. We also plan to extend our deterministic framework to the stochastic setting, as population heterogeneity is known to further complicate the control process. This work demonstrates the benefits of combining OC and RL to simplify the control problem, showing how their frameworks can be effectively integrated. This connection can be further developed by using the memoryless OC solution to enhance RL training through reward shaping (Ng et al., 1999; Wiewiora, 2003; Adamczyk et al., 2023) or pre-training (Uchendu et al., 2023) techniques.
In this work we present a strong proof of concept for RL-guided drug dosing, which we hope will motivate further experimental work on adaptive therapies. Since preparing drug doses of arbitrary concentration can be challenging, the derived bang-bang control simplifies the practical implementation of our policies. While our results show that the optimality of such policies is robust against memory strength, switching rates, and dose-response curves, their optimality in the clinical setting may be influenced by additional factors such as patient well-being, interactions with the host system, and penalties for drug exposure. Addressing these constraints highlights the need to bridge the “sim-to-real” gap through both novel computational approaches and evaluations on real-world data.
Inspired by its success in robotics applications (Akkaya et al., 2019), domain randomization may prove useful in addressing such computational challenges. By training agents in a diverse set of simulated environments with randomized parameters (e.g., switching rates, dose-response effects, noise levels), policies can become more robust for real-world application. With such general agents successfully trained, an offline RL problem (Levine et al., 2020) with warmstarting (Uchendu et al., 2023; Agarwal et al., 2022) could be formulated once real-world data is available. We hope that this approach can provide a path toward generalization, thus enabling more reliable application of RL policies in diverse real-world contexts.
8 Discussion
In this work, we study the control of a highly non-Markovian model of adaptive cellular growth dynamics using deep reinforcement learning. Although we focus here on a specific parameterization most relevant to cancer, we expect this methodology to be applied successfully to other scenarios, including resistance development in bacteria. We find that deep RL is capable of recovering the known optimal policy for the memoryless case and can successfully find a policy for memory-based systems which prevents proliferation, all without access to the underlying model parameters. Furthermore, we obtain a policy that is robust to measurement noise and memory strength perturbations.
Exact methods from optimal control require direct access to the population’s resistant fraction and underlying model parameters to determine switching rates. In practice, experimentally measuring the growth rate of a pathogenic population is simpler than determining its corresponding resistant fraction. Surprisingly, we find that model-free reinforcement learning can provide successful policies directly from the growth rate itself, making it a promising method for use in clinical settings.
9 Acknowledgements
JA acknowledges funding support the NSF through Award No. PHY-2425180; the use of the supercomputing facilities managed by the Research Computing Department at UMass Boston; and fruitful discussions with Rahul V. Kulkarni and Stas Tiomkin. JCK acknowledges support from the Computational Biology Department at CMU, and from Shiladitya Banerjee in the Department of Physics at CMU. This project developed out of the hackathon at the 2024 Summer School hosted by the Institute for Artificial Intelligence and Fundamental Interactions (IAIFI).
Appendix A Reinforcement Learning Details
@@
We adapted DQN from Stable-Baselines3 (Raffin et al., 2021) with a Double DQN action selection rule (Van Hasselt et al., 2016). We train the RL agent for $3\times 10^{5}$ total environment steps. We limited the episodes to be of length $10^{4}$ steps (corresponding to 100 hours in simulation). An MLP of fixed size ($2$ hidden layers with $64$ dimensions each and ReLU activation) was used to parameterize the $Q$-function.
@@
Regarding the environment and training, we list several key implementation choices: At the beginning of an episode, the state is zero-padded to always be of length $K=5$. We experimented with adding a penalty for allowing the number of cells to increase beyond the initial amount upon termination, but found this was not necessary for successful training. Borrowing terminology from the literature on Atari environments (Mnih et al., 2015), we stack $K$ frames (previous cost values) to form the RL agent’s state vector. Although initially we let $K$ be a $\mu$-dependent hyperparameter, we found a constant choice of $K=5$ to work well across the values of $\mu$ studied.
@@
A.1 Hyperparameters
We find that sweeping over a range of hyperparameters (as shown in Table A.1) did not have a significant effect on performance, though for reproducibility we list the final hyperparameters used (for $\mu=0.7$) below for DQN and FQF experiments, respectively. (N/A indicates the value was not swept over, and fixed to the finetuned value.) The discount factor $\gamma=0.999$, gradient steps per update ($1$), frames stacked ($5$) and buffer size ($100{,}000$) are fixed throughout.
Table 2: Hyperparameters for Double DQN
Hyperparameter Sweep Values Finetuned Value
batch size $16$, $32$, $64$ $32$
exploration rate $0.01-0.2$ $0.05$
learning rate $10^{-5}-10^{-3}$ $3.60\times 10^{-4}$
target update interval $1{,}000,5{,}000,10{,}000,30{,}000$ $1{,}000$
Table 3: Hyperparameters for (NoisyNet) FQF
Hyperparameter Sweep Values Finetuned Value
batch size $8$, $64$ $8$
exploration rate $0.02-0.2$ $0.2$
learning rate $5\times 10^{-4}-5\times 10^{-3}$ $2.7\times 10^{-3}$
target Polyak averaging $0.001,0.005,0.01$ $0.005$
$n$-step TD $1,2,4,32$ $1$
entropy coeff. $0,0.0005,0.001,0.003$ $0.003$
Appendix B Further Experimental Results
Figure 6: We find that the simulation time can have a significant effect on RL performance. For each choice of $\Delta$, we run a random sweep of size $30$ over various hyperparameters, selecting the highest-performing run for each $\Delta$. Since smaller values of $\Delta$ require longer compute-times for simulations, there is a tradeoff between the amount of time (also, the maximum pulsing frequency, relevant in clinical settings) and the best performance. We have chosen to use $\Delta=0.01$ throughout.
Figure 7: Robustness experiment on changing switching rates $\alpha,\ \delta$. The black box in the center highlights the rates at which the agent was trained. On each axis, the rate is varied by ±$\pm 20\%$, representing a large range of biologically plausible parameters. As $\alpha$ increases (or $\delta$ decreases), the cells are more often in the resistant state (cf. Fig. 1), making the control problem more difficult for the agent.